From the information, determine the ratio of pressures. Hvap of 32.1 kJ/mol, 35°C and 117°C
options:
A) 7.29×10^32
B) 13.8
C) cannot calculate
D) 0.072
E) 1.37×10^-33
Answer #1

According to Clapeyron equation :

ln (P2/P1) = -\DeltaHvap/R (1/T2-1/T1)

T1 = 35 +273.15 = 308.15 K

T2 = 117 +273.15 = 390.15 K

\DeltaHvap = 32.1 kJ/mol = 32.1 x 1000 = 32100 J/mol

R = 8.314 J/mol K

ln (P2/P1) = - 32100/8.314 ((1/390.15)-(1/308.15))

ln (P2/P1) = -3860.96 ((0.00256 - 0.00324))

ln (P2/P1) = -3860.96 ((-0.00068)

ln (P2/P1) = 2.625

(P2/P1) = e2.625 = 13.80

P2/P1 = 13.80

P1/P2 = 1/13.80 = 0.072